# Hybrid Unsupervised Learning Strategy for Monitoring Industrial Batch Processes

### v1

@@ 4.5 Anomaly detection
Anomaly detection is a crucial aspect of a comprehensive monitoring system and involves identifying atypical patterns, events, or behaviors in dataset that significantly deviate from normal behaviour. For the comparative evaluation of the anomaly detection performance, another batch sequence was recorded. As shown in fig. 4(a) the sequence contains six batches, of which the three sequences E1, E2 and E3 deviate from the normal behavior. In batch E1 the venting of reservoir B02 (cf. fig. 1) has been reduced, in batch E2 the flow cross-section has been decreased by valve V02 and in batch E3 the level gauge position in tank B01 has been shifted. Batches N3 to N5 reflect the normal behavior of the process. In fig. 4 (b) and (c) the time courses of the Euclidean distance $e_{i\mathbf{v^{*}}}$ as well as the identified process phases $c_{i}$ are visualized. Focusing on batches N3, N4, and N5, where normal behavior is expected, shows that the standard SOM model is very sensitive to new, unseen data, resulting in a high $e_{i\mathbf{v^{*}}}$. This phenomenon, already described in section 4.3, is problematic with respect to a real industrial application in the sense that false alarms can be generated (e.g. batch N4). In terms of anomaly detection performance, both algorithms perform very well in terms of reliably detecting the anomalies in the batches E1, E2, and E3, but the HULS procedure still performs better in terms of recovering from the error state. Finally, regarding the process phase assignment $c_{i}$, it is clear that the standard SOM model cannot correctly identify the process phase because, as noted in section  2, only three phases could be identified.
Figure 4: (a) Recorded Batch sequence for evaluating anomaly detection performance. (b) (b) Resulting quantification error $e_{i\mathbf{v^{*}}}$ and class assignment $c_{i}$ for standard SOM model and (c) for proposed HULS procedure.

### v2

@@ 4.5 Anomaly detection
Anomaly detection is a crucial aspect of a comprehensive monitoring system and involves identifying atypical patterns, events, or behaviors that significantly deviate from normal behaviour. For the comparative evaluation of the anomaly detection performance, another batch sequence was recorded. As shown in Fig. 4(a) the sequence contains six batches, of which the three sequences E1, E2 and E3 deviate from the normal behavior. In batch E1 the venting of reservoir B02 (cf. Fig. 1) has been reduced, in batch E2 the flow cross-section has been decreased by valve V02 and in batch E3 the level gauge position in tank B01 has been shifted. Batches N3 to N5 reflect the normal behavior of the process. In Fig. 4 (b) and (c) the time courses of the Euclidean distance $e_{i\mathbf{v^{*}}}$ as well as the identified process phases $c_{i}$ are visualized. Focusing on batches N3, N4, and N5, where normal behavior is expected, shows that the standard SOM model is very sensitive to new, unseen data, resulting in a high $e_{i\mathbf{v^{*}}}$. This phenomenon, already described in section 4.3, is problematic with respect to a real industrial application in the sense that false alarms can be generated (e.g. batch N4).
In terms of anomaly detection performance, both algorithms perform very well in terms of reliably detecting the anomalies in the batches E1, E2, and E3, but the HULS procedure still performs better in terms of recovering from the error state. Finally, regarding the process phase assignment $c_{i}$, it is clear that the standard SOM model cannot correctly identify the process phase because, as noted in section  2, only three phases could be identified.
Figure 4: (a) Recorded Batch sequence for evaluating anomaly detection performance. (b) Resulting quantification error $e_{i\mathbf{v^{*}}}$ and class assignment $c_{i}$ for standard SOM model and (c) for proposed HULS procedure.